मराठी
तामिळनाडू बोर्ड ऑफ सेकेंडरी एज्युकेशनएचएससी विज्ञान इयत्ता १२

The probability density function of the random variable X is given by ef(x)={16xe-4xx>00x≤0Find the mean and variance of X

Advertisements
Advertisements

प्रश्न

The probability density function of the random variable X is given by

`f(x) = {{:(16x"e"^(-4x), x > 0),(0, x ≤ 0):}`
find the mean and variance of X

बेरीज
Advertisements

उत्तर

Given p.d.f is `f(x) = {{:(16x"e"^(-4x), x > 0),(0, x ≤ 0):}`

Mean: E(X)

= `int_-oo^oo x f(x) "d"x`

= `16int_0^oo x^2 "e"^(-4x) "d"x`

 Using integration by parts method twice

Let u = x2

⇒ du = 2x dx

And `int "dv" = int "e"^(-4x)`

v = `"e"^(-4x)/(-4)`

`int "u" "d" = "uv" - int "v"  "du"`

`int^2"e"^(-4x) "d"x = (x^2"e"^(-4x))/(-4) + 1/4 int 2x"e"^(-4x) "d"x`  .......(1)

= `- (x^2"e"^(-4x))/4 + 1/2 int x"e"^(-4x) "d"x`

∵ Integration by parts method

u = x

⇒ du = dx

And `int "dv" - int"e"^(-4x) ""x`

v = `"e"^(-4x)/(-4)`

`int "u"  "dv" = "uv" - int "v"  "du"`

`int x"e"^(-4x) "d"x = (-x"e"^(-4x))/4 + 1/4 int "e"^(-4x) "d"x`

= `(-x"e"^(-4x))/4 - 1/16 "e"^(-4x)`

Substituting in (1)

`intx^2"e"^(-4x) "d"x = (x^2"e"^(-4x))/(-4) + 1/2 [(-x"e"^(-4x))/4 - 1/16 e"^(-4x)]`

E(X) = `16[(x^2"e"^(-4x))/(-4) - (x"e"^(-4x))/8 - "e"^(-4x)/32]_0^oo`

= `16[0 - ((-1)/32)]`

= `16[1/32]`

= `1/2`

E(X2] = `int_-oo^oo x^2 f(x) "d"x`

= `16int_0^oo x^3"e"^(-4x)  "d"x`

Using integration by parts method

Let u = x3

⇒ du = 3x2 du

And `int "dv" = int "e"^(-4x) "d"x`

⇒ v = `"e"^(-4x)/(-4)`

`int "u" "dv" = "uv" - int "v" "du"`

`intx^3"e"^(-4x) "d"x = - (x^3"e"^(-4x))/4 + 3/4 int"e"^(-4x) x^2 "d"x`

= `- (x^3"e"^(-4x))/4 + 3/4[(x^2"e"^(-4x))/(-4) - (x"e"^(-4x))/8 - "e"^(-4x)/32]`

∵ Using E(X) integration]

= `- (x^3"e"^(-4x))/4 - 3/16 x^2"e"^(-4x) - 3/32 x"e"^(-4x) - 3/128 "e"^(-4x)`

E(X2) = `16[- (x^3""^(-4x))/4 - 3/16 x^2"e"^(-4x) - 3/32 x"e"^(-4x) - 3/128 "e"^(-4x)]_0^oo`

= `16[0 - ((-3)/128)]`

= `16[3/128]`

= `3/8`

Variance Var(X) = E(X2) – [E(X)]2 

= `3/8 - 1/4`

= `(3 - 2)/8`

= `1/8`

shaalaa.com
Mathematical Expectation
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 11: Probability Distributions - Exercise 11.4 [पृष्ठ २१०]

APPEARS IN

सामाचीर कलवी Mathematics - Volume 1 and 2 [English] Class 12 TN Board
पाठ 11 Probability Distributions
Exercise 11.4 | Q 7 | पृष्ठ २१०

संबंधित प्रश्‍न

For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(2(x - 1), 1 < x ≤ 2),(0, "otherwise"):}`


Two balls are drawn in succession without replacement from an urn containing four red balls and three black balls. Let X be the possible outcomes drawing red balls. Find the probability mass function and mean for X


Four fair coins are tossed once. Find the probability mass function, mean and variance for a number of heads that occurred


A commuter train arrives punctually at a station every half hour. Each morning, a student leaves his house to the train station. Let X denote the amount of time, in minutes, that the student waits for the train from the time he reaches the train station. It is known that the pdf of X is

`f(x) = {{:(1/30, 0 < x < 30),(0, "elsewhere"):}`
Obtain and interpret the expected value of the random variable X


Choose the correct alternative:

Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he loses ₹ k2, where k is the face that comes up k = {1, 2, 3, 4, 5}. The expected amount to win at this game in ₹ is


Choose the correct alternative:

Four buses carrying 160 students from the same school arrive at a football stadium. The buses carry, respectively, 42, 36, 34, and 48 students. One of the students is randomly selected. Let X denote the number of students that were on the bus carrying the randomly selected student. One of the 4 bus drivers is also randomly selected. Let Y denote the number of students on that bus. Then E(X) and E(Y) respectively are


Choose the correct alternative:

On a multiple-choice exam with 3 possible destructive for each of the 5 questions, the probability that a student will get 4 or more correct answers just by guessing is


Choose the correct alternative:

If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is


Let X be a continuous random variable with probability density function
`"f"_x(x) = {{:(2x",", 0 ≤ x ≤ 1),(0",",  "otherwise"):}`
Find the expected value of X


In investment, a man can make a profit of ₹ 5,000 with a probability of 0.62 or a loss of ₹ 8,000 with a probability of 0.38. Find the expected gain


Let X be a random variable and Y = 2X + 1. What is the variance of Y if variance of X is 5?


Choose the correct alternative:

Demand of products per day for three days are 21, 19, 22 units and their respective probabilities are 0.29, 0.40, 0.35. Profit per unit is 0.50 paisa then expected profits for three days are


Choose the correct alternative:

Given E(X) = 5 and E(Y) = – 2, then E(X – Y) is


Choose the correct alternative:

If X is a discrete random variable and p(x) is the probability of X, then the expected value of this random variable is equal to


Choose the correct alternative:

Which of the following is not possible in probability distribution?


Prove that V(aX) = a2V(X)


Prove that V(X + b) = V(X)


The time to failure in thousands of hours of an important piece of electronic equipment used in a manufactured DVD player has the density function
f(x) = `{{:(2"e"^(-2x)",",  x > 0),(0",",  "otherwise"):}`
Find the expected life of this piece of equipment


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×